# Electron-mediated anharmonicity and its role in the Raman spectrum of graphene

### v1

@@ Introduction
Out-of-equilibrium conditions. We now turn to a slightly different scenario achieved in ultrafast Raman experiments. The main distinction of the ultrafast regime with respect to the continuous wave measurements is that thermal equilibrium between the electron and phonon subsystems is not achieved Yan et al. (2009); Ferrante et al. (2018). In one particular case, where , we expect the EPC to play the leading role, as anharmonic effects are not driven by the electronic temperature. We employ the same methodology as for the equilibrium case, but instead of setting the G mode temperature equal to the electronic one, we keep the G mode at 300 K, and vary the electronic temperature Te. In order to simulate the experiment done in Ref. 13, we perform our calculations on doped graphene with meV. In Fig. 7 we show the inverse lifetime of electron-hole pairs and their energy renormalization, calculated by progressively more complete approaches. As for the equilibrium case, taking into account the temperature-dependent quasiparticle DOS leads to the largest contributions to both quantities. However, for , an order of magnitude is lost in comparison with respect to results presented in Fig. 3, due to the large magnitude of the Bose-Einstein distribution factors for high phonon temperatures. The G mode frequency dependence in the non-equilibrium regime is now almost entirely the result of the interband phonon self-energy. The experimentally observed trend is reversed from the equilibrium one (i.e., increase of the frequency with increasing temperature) Yan et al. (2009); Ishioka et al. (2008); Ferrante et al. (2018). Our calculations explain this is due to the decreased impact of the anharmonic and EPC intraband contribution, and reproduce the slope very well. The temperature dependence of the linewidth does not change qualitatively in comparison with the equilibrium case, but its magnitude is reduced. The key role belongs to the intraband EPC contribution. We consider EPC to be the driving mechanism behind the temperature dependence in the nonequilibrium case, making the accurate calculation of the EPC phonon self-energy even more essential.
@@
In this work, we revisit the EPC theory in graphene and reveal its effect on the temperature-dependent properties of the G mode Raman spectral features. The long-wavelength phonons are screened by the first-order interband NA coupling and the higher-order dynamical intraband term, which can be seen as electron-mediated phonon-phonon coupling. The latter effect is essentially anharmonic and, in combination with conventional anharmonicity, captures the full temperature dependence of G mode frequency and linewidth, in line with experimental results. We believe this type of approach is crucial to describe Raman experiments for materials with significant coupling of electrons to Raman active modes. In graphene we find that the EPC-induced and conventional anharmonicities play different roles in different doping regimes. In undoped samples, standard anharmoncity is the dominant contribution to the phonon dynamics, while for the moderately- to highly-doped graphene, the electron-mediated phonon-phonon coupling is crucial. The latter EPC effect is also essential to understand Raman features in non-equilibrium graphene. As for the lattice-driven anharmonicity, by comparing SSCHA and TDEP calculations, we show that four-phonon scattering contributions are more important than the three-phonon scattering for the G phonon linewidth.
@@
In conclusion, we emphasize that the present theory could be applied to any material with strong non-adiabatic effects Girotto and Novko (2023a), such as doped layered materials like transition metal dichalcogenides 9, hole-doped diamond Caruso et al. (2017), or MgB2-related compounds Yu et al. (2022). The present EPC effects might also explain the non-monotonous linewidth behavior in Weyl semimetals, as anharmonic effects are unable to describe the observed temperature dependence Coulter et al. (2019); Wulferding et al. (2020); Osterhoudt et al. (2021); Yu et al. (2023); Cheng et al. (2024). In some charge-density-wave materials, it could be the leading mechanism behind the CDW transition Varma and Simons (1983) and it could also be important for strongly coupled or soft phonons away from the Brillouin zone center Flicker and van Wezel (2015). Even more, our EPC theory could explain the non-monotonous temperature dependence of the phonon linewidth observed for the pristine Ru(0001) surface Liu et al. (2024).
@@ Methods
Electron-phonon contribution. From first principles we calculate electron and phonon properties with QUANTUM ESPRESSO Giannozzi et al. (2009, 2017, 2020) and EPC with the EPW code Giustino et al. (2007); Noffsinger et al. (2010); Poncé et al. (2016). We use norm-conserving scalar pseudopotential with an energy cutoff of 100 Ry. The relaxed lattice constant is and the periodic graphene planes are separated by . We use the Fermi-Dirac smearing with K.

### v2

@@ Results
Out-of-equilibrium conditions. We now turn to a slightly different scenario achieved in ultrafast Raman experiments. The main distinction of the ultrafast regime with respect to the continuous wave measurements is that thermal equilibrium between the electron and phonon subsystems is not achieved Yan et al. (2009); Ferrante et al. (2018). In one particular case, where , we expect the EPC to play the leading role, as anharmonic effects are not driven by the electronic temperature. We employ the same methodology as for the equilibrium case, but instead of setting the G mode temperature equal to the electronic one, we keep the G mode at 300 K, and vary the electronic temperature Te. In order to simulate the experiment done in Ref. 13, we perform our calculations on doped graphene with meV. In Fig. 7 we show the inverse lifetime of electron-hole pairs and their energy renormalization, calculated by progressively more complete approaches. As for the equilibrium case, taking into account the temperature-dependent quasiparticle DOS leads to the largest contributions to both quantities. However, for , an order of magnitude is lost in comparison with respect to results presented in Fig. 3, due to the large magnitude of the Bose-Einstein distribution factors for high phonon temperatures. The G mode frequency dependence in the non-equilibrium regime is now almost entirely the result of the interband phonon self-energy. The experimentally observed trend is reversed from the equilibrium one (i.e., increase of the frequency with increasing temperature) Yan et al. (2009); Ishioka et al. (2008); Ferrante et al. (2018). Our calculations explain that this is due to the decreased impact of the anharmonic and EPC intraband contribution, and reproduce the slope very well. The temperature dependence of the linewidth does not change qualitatively in comparison with the equilibrium case, but its magnitude is reduced. The key role belongs to the intraband EPC contribution. We consider EPC to be the driving mechanism behind the temperature dependence in the nonequilibrium case, making the accurate calculation of the EPC phonon self-energy even more essential.
Discussion
@@
In this work, we revisit the EPC theory in graphene and reveal its effect on the temperature-dependent properties of the G mode Raman spectral features. The long-wavelength phonons are screened by the first-order interband NA coupling and the higher-order dynamical intraband term, which can be seen as electron-mediated phonon-phonon coupling. The latter effect is essentially anharmonic and, in combination with conventional anharmonicity, captures the full temperature dependence of G mode frequency and linewidth, in line with experimental results. We believe this type of approach is crucial to describe Raman experiments for materials with significant coupling of electrons to Raman active modes. In graphene we find that the EPC-induced and conventional anharmonicities play different roles in different doping regimes. In undoped samples, standard anharmoncity is the dominant contribution to the phonon dynamics, while for the moderately- to highly-doped graphene, the electron-mediated phonon-phonon coupling is crucial. The latter EPC effect is also essential to understand Raman features in non-equilibrium graphene. As for the lattice-driven anharmonicity, by comparing SSCHA and TDEP calculations, we show that four-phonon scattering contributions are more important than the three-phonon scattering for the G phonon linewidth, and that the quantum thermal expansion effects are crucial for obtaining the right values for the four-phonon part.
@@
In conclusion, we emphasize that the present theory could be applied to any material with strong non-adiabatic effects Girotto and Novko (2023a), such as doped layered materials like transition metal dichalcogenides 9, hole-doped diamond Caruso et al. (2017), or MgB2-related compounds Yu et al. (2022). The present EPC effects might also explain the non-monotonous linewidth behavior in Weyl semimetals, as the standard anharmonic effects are unable to describe the observed temperature dependence Coulter et al. (2019); Wulferding et al. (2020); Osterhoudt et al. (2021); Yu et al. (2023); Cheng et al. (2024). In some CDW materials, it could be the leading mechanism behind the CDW transition Varma and Simons (1983) and it could also be important for strongly coupled or soft phonons away from the Brillouin zone center Flicker and van Wezel (2015). In that case, it could provide a leading contribution to the formation of the gap or pseudogap Yoshiyama et al. (1986), as well as to the value of the CDW transition temperature Varma and Simons (1983). Furthermore, our EPC theory could explain the non-monotonous temperature dependence of the phonon linewidth observed for the pristine Ru(0001) surface Liu et al. (2024).
@@ Methods
Electron-phonon contribution. From first principles we calculate electron and phonon properties with QUANTUM ESPRESSO Giannozzi et al. (2009, 2017, 2020) and EPC with the EPW code Giustino et al. (2007); Noffsinger et al. (2010); Poncé et al. (2016). We use a norm-conserving scalar pseudopotential from PseudoDojo van Setten et al. (2018) with an energy cutoff of 100 Ry. The relaxed lattice constant is and the periodic graphene planes are separated by . We use the Fermi-Dirac smearing with K.